OpenStudy (anonymous):

The sum of four consecutive odd integers is 28458968. What is the second number? Any help pliz how to work out ???

OpenStudy (anonymous):

An odd number can be expressed as 2k+1 for some integer k. So the sum of four consecutive odd numbers can be expressed as \[((2k + 1) - 4) + ((2k + 1) - 2) + (2k+1) + ((2k + 1) + 2)\]\[= (2k-3) + (2k-1) + (2k +1) + (2k+3)\] Try to use that expression to get the answer.

OpenStudy (anonymous):

As fruitbasket said, any odd number can be expressed as 2k+1 for any integer k. So four consecutive odd numbers are 2k+1, 2k+3, 2k+5 and 2k+7 and their sum is 8k+16. Therefore, 8k+16=28458968 and k=3557369

OpenStudy (anonymous):

thanks mate for the helpful answer

OpenStudy (anonymous):

Just make sure that you substitute k into whatever expression you used for the 2nd odd number to get its value.

OpenStudy (anonymous):

expression above can work for both even and odd consecutive integers??

OpenStudy (anonymous):

2k is always an even integer (as long as k is an integer) due to the factor of 2. Odd numbers are just even numbers with 1 added to them, so: CONSECUTIVE EVEN NUMBERS: \[2k + (2k + 2) + (2k + 4) + ...\]CONSECUTIVE ODD NUMBERS: \[(2k+1) + (2k+3) + (2k+5)+...\]

OpenStudy (agent0smith):

If the first odd integer is x, then the second (the one you want to find) is x+2, the third is x+4, then the fourth is x+6. Since the sum is 28458968, add them all together: \[\large x + (x+2) + (x+4) + (x+6) = 28458968 \]

OpenStudy (anonymous):

thanks very much for your help